Libby Area Technical Assistance Group, Inc.
PO Box 53 • Libby, MT 59923
November 6th, 2007 • Regular Board Meeting Minutes
(Note: bold items within paragraphs are motions made and voted on as well as action items agreed upon.)
Board Member Attendees:
- Mike Noble
- Helen Andries
- JoElyn Brus
- LeRoy Thom
- Eileen Carney
- Alvin Benitz
- Dr. Brad Black
- Nancy Hogan
- Cheryl Fox
- Armstrong, Ward
- Budeau, D.
- Cirian, Mike
- Cramer, Darin
- Erickson, Don
- Erquiaga, Phillip
- Fitzsimmons, Dave
- Harrison, Marla
- Kaplowe, Dave
- Kucos, Steve
- LaCroix, Mark
- LeCours, Catherine
- Linnert, Ted
- Magenheimer, Sara
- Mannix, R.
- McKay, Toby
- Moore, Sara
- O’Ferrall, Fred
- Paul, Jim
- Pope, Michael
- Pope, Nagle
- Ramos, Pedro
- Ries, Bob
- Ross, Joe
- Rykoff, Andri
- Sullivan, Gordon
- Ventra, Carrie
- Watts, Lee
- Wolfer, B.
- Young, Bob
- Zimmerman, Scott
1. Meeting began at approximately 7:00pm
3. Meeting minutes to be reviewed:
a. Meeting minutes from the October 9th, 2007 regular board meeting were reviewed.
i. Eileen Carney made the motion to accept the meeting minutes as written. JoElyn Brus second the motion. Motion unanimously passed.
b. The Technical Advisor contract has been reviewed and presented to the EBoard.
i. Eileen Carney made the motion to accept the contract and forward to EPA for review. Helen Andries second the motion. Motion unanimously passed.
c. The Grant Administrator contract was reviewed and presented to the EBoard.
i. Helen Andries made the motion to accept the contract and forward to EPA for review. JoElyn Brus second the motion. Motion unanimously passed.
4. Treasurer’s report – JoElyn Brus
a. Current grant balance reported is $40,190.58, this does not reflect a pending reimbursement (#49) in the amount of $682.56.
5. Old Business
a. Additional information is needed regarding the environmental assessment and the due date of this document. A request has been received for quarterly updates. (ACTION ITEM).
6. EPA Report (work in progress)
a. Libby Area – Mike Cirian
i. EPA has finished for the season, one property was scheduled for partial demolition, and is currently being fully demolished this week (702 Main Street).
ii. For the year 2007 it was estimated that 160 properties would be done, and 160 properties were cleaned. ERS activities have cut into the clean up budget.
1. January of 2008 a report will be available regarding the year 2007.
iii. Project total of properties cleaned is 954. Approximately 45,000 cubic yards of soil have been removed this year. This equates to just shy of a total of 500,000 cubic yards of soil removed for the project.
1. Current average clean up size is just over three acres.
2. A total of 1,650 cubic yards of VCI has been moved.
iv. Summary – Where do we go from here?
1. Issues will be addressed regarding how the 2007 cleanup occurred and items will be reviewed regarding homeowners and businesses, improvements on containment and construction, breach of protocols and general health and safety issues.
2. Mine decontamination is an area of needed improvement, an outside agency for improvements will be used.
1. LeRoy Thom asked if there were any penalties for noncompliance. RESPONSE: There is a cure notice sent with a ten day correction plan and time frame required. A default could occur which would result in non-payment to the contractor who is out of compliance. Non-compliance could impact the contractor being awarded future government contracts.
2. Gordon Sullivan asked to have the mine site outlined. RESPONSE: There is an amphitheatre type area which is the actual deposit of contaminated material. Equipment is isolated at the mine site until the end of year. Water is pumped from the pond for decontamination and the decontaminated waters are pumped into the amphitheatre. Mike Cirian believes that the water drains from the amphitheatre back to the pond. At the end of the year equipment is decontaminated and returned to town/storage. Only equipment/vehicles with positive pressure units or full tyvec is used at the mine site..
3. It was mentioned that pollution from the Rainey Creek comes from the mine site and other forest service land in the area. Catherine LeCours said that Bonnie Lavel is working the issues and classifications of streams. They are studying the load allocations, source of pollution and possible responsible party. The goal is to meet surface and ground water regulations.
a. EPA is working with the State of Montana to meet regulations, they are just starting on the mine site.
vi. There will be a meeting in December at EPA (in Denver) with the contractors regarding the events in 2007. Discussion will be on financial needs for 2008 and what will happen with the continued cleanup.
b. Troy Area: (Catherine LeCours)
i. They have screened 536 properties, and their goal was 450.
ii. They are working with the State of Montana on an agreement with EPA for the Troy cleanup. They are currently evaluation 1,200 properties with about 30% of the evaluated properties needing removal based on visual standards.
iii. They are utilizing the emergency removal criteria.
1. Gordon Sullivan asked if they are utilizing the ERS contact line. RESPONSE: Yes, one ERS call was from a new homeowner allowing access (previous homeowner refused access), and another call was in regarding to remodeling. Approximately 300 to 500 properties remain to be screened.
iv. Property access agreements are being received and currently for year 2008 they have six high priority cleanup properties. These properties have a higher priority than Libby properties. They are working with ERS on containment issues.
v. They are working with the inspections (screening), discussions with homeowners, they will be conducting a mass mailing in February or March of 2008.
vi. Challenges exist with 160 properties cleaned a year giving an average of just over three years remain for the cleanup. Also gaining access to properties continues to be a challenge, in addition to the changing of ownership of properties. Prioritizing properties will also be a challenge.
1. Alvin Benitz asked about the budget for the Troy cleanup. RESPONSE: There is no budget at this time.
2. Other challenges discussed is the weigh scale has been removed for hauling loads, where will be waste be hauled to (possibly Libby), or will the Troy landfill be expected to take the contaminated material. What is the volume of the waste material expected? Troy hopes for a quick and efficient way to get the cleanup done. There will be public hearing if the landfill will be used.
3. At this time the percentage of properties needing either exterior or interior cleanup is unknown.
4. About 15 to 20 properties are refusing access for the cleanup.
5. Comments were made regarding the real-estate disclosure laws require properties for sale to disclose the possibility of vermiculite contamination. Dr. Brad Black commented that the LATAG newsletter could help in the areas of education. They need to be distributed at the drug stores and hardware stores in both Libby and Troy. The Troy office can distribute some newsletters; they are open Tuesday, Wednesdays and Thursdays.
7. Technical Advisor Report – Dr. Gerry Henningsen
a. No Report - work stoppage
8. Sub-Committee Reports
a. Communications – Helen Andries (chairperson)
i. Web site – it was reported that LATAG is not linked to EPA and Ted Linnert offered his help with any web links.
ii. The newsletter is still pending.
b. ERS – Mike Noble (chairperson) [Reported by Mike Cirian]
i. Not many calls were received; one involved a remodel of a bathroom.
ii. Advanced notice is needed – emergency responses are very costly.
1. Gordon Sullivan asked when the ERS position was moving out from EPA. RESPONSE: A ROD is needed then O&M will most probably address these issues.
c. Nomination – Helen Andries (chairperson)
i. LATAG received a nomination form from Phillip Erquiaga. Helen Andries made the motion to accept Phillip as a new board member, JoElyn Brus second the motion. Motion unanimously passed.
ii. The floor was opened for nominations for the Ad-Hoc position that will need to be replaced in January of 2008.
1. JoElyn Brus nominated LeRoy Thom to continue holding the Ad-Hoc position.
2. Eileen Carney nominated LeRoy Thom to continue holding the Ad-Hoc position.
3. Nominations closed
d. O&M – Mike Noble/LeRoy Thom (co-chairpersons)
i. No LATAG attendance at the last meeting. Catherine LeCours reported that they have started a working document regarding the maintenance of the export plant. Paul Lammers has been tasked to put the plan on paper. EPA has copies available to the public.
e. Technical Advisory – Dr. Brad Black (chairperson)
i. Epidemiology work is moving forward. Work is progressing with EPA on the database with Dr. Steve Levin’s involvement. They would like to study a population sub-set with low exposure and follow individuals over time for any early signs of symptoms to low exposures.
ii. Toxicology studies are being done but no status of these studies is available yet.
9. New Business (none)
10. Special Guests – Ward Armstrong – Portland State University
a. A group of visits from the state of Washington are here to learn about the superfund and the processes Libby is going through regarding the cleanup. They are studying what is being done and how it might be applied to other agencies.
b. They are providing mid-career leadership training regarding public policy issues as related to national resources.
c. They want to know how public policy decisions are made and how to improve on these decisions. This involves all issues involving government decisions. Because Libby was declared a Superfund site they are visiting our area.
11. Open Discussion
12. Next Meeting Scheduled December 11th, 2007.
a. Helen Andries made the motion to adjourn, JoElyn Brus second the motion. Motion unanimously passed.
Meeting adjourned at approximately 8:30pm